Hi
There was enough food for 72 men to last 24 days. After 4 days another 8 men join them. For how many days will the remaining food last.
Thanks

Answer by greenestamps(13195)



In the beginning, there was enough food for 72 men for 24 days.

After 4 days, there was enough food remaining for 72 men for 20 days.

When 8 more men joined the group, there were now 80 men; the number of men increased by a factor of 80/72 = 10/9.

That means the number of days the food would last DECREASES by a factor of 10/9.

ANSWER: 20*(9/10) = 18 days